OpenStudy (anonymous):

solve the equation ((7t+14)/(2t2+5t−12))+(1/(t+4))=(4/(2t−3)) i already know the answer i just need to know how to get it

OpenStudy (anonymous):

opps had a mistake let me repost it

OpenStudy (anonymous):

putz!

OpenStudy (anonymous):

((7t+14)/(2t^2+5t−12))+(1/(t+4))=(4/(2t−3))

OpenStudy (anonymous):

i was considered that! =)

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

i found t= 9/12 its correct? If it is, i wrote how i found..

OpenStudy (anonymous):

i got t=1 but i cant remember how i got it and i gotta practice for mt test.

OpenStudy (anonymous):

im wrong!

OpenStudy (anonymous):

-_- i wish i could remember...

OpenStudy (anonymous):

\o/

OpenStudy (anonymous):

i was forgetting the 2 on first denominator!

OpenStudy (anonymous):

you can rewrite the first denominaotr as 2(t-1.5)(t+4) so multiplying both sides by 2(t-1.5)(t+4) and solving to t we get t=1!

OpenStudy (anonymous):

ah! we need to rewrite (2t-3) as 2(t-1.5)

OpenStudy (anonymous):

(7t+14) +2(t-1.5) -4(t+4)= 0 t=1

OpenStudy (anonymous):

i got it 5t-5=0 thus making t=1

OpenStudy (anonymous):

thanks though!

OpenStudy (anonymous):

you're welcome